OpenStudy (anonymous):

Need some help learning how to use Elimination to solve a system of three equations for X, Y, and Z. I am supposed to enter the numbers like an ordered triple, like this, (2, -3, 4) Here I have these... { 4x + y + 3z = 0 { 2x - 2y - z = 10 { 3x - 2y + 2z = 11

OpenStudy (anonymous):

I am unsure of how to go about this. What should I do first?

OpenStudy (wilsondanielle):

Isolate one of the variables and use that to find the values of the other. I would start with the first equation because it is equal to 0 and will be more simple. For example, if you were isolating for x, you would end up with 4x = y+3z x=0.25(y=3z) You could then substitute that value of x into one of the other equations, and then you would develop it and isolate for another variable in the same fashion. Using the third equation, you would find a true value of the remaining variable and then use that to find the values of the other two.

OpenStudy (wilsondanielle):

*0.25(y+3z)